OpenStudy (anonymous):

is the sqrt of 4 - (-5) undefined?

OpenStudy (anonymous):

because you cant have a negative for a sqrt?

OpenStudy (ranga):

when you simplify 4 - (-5) what do you get?

OpenStudy (aakashsudhakar):

√(4 - -5) √(4 + 5) √9 The square root of (4 - (-5)) is defined, because even though you can't take the sq.rt. of a negative number, the quantity isn't a negative number. 4 - (-5) is the same as saying 4 + 5, which just equals 9. You can take the square root of 9.

OpenStudy (anonymous):

ohhhh 9

OpenStudy (anonymous):

you can. at first i also thought it was sqrt(4 -5) = sqrt(-1). but it's not

OpenStudy (dan815):

ya silly

OpenStudy (anonymous):

okay thnx =X

OpenStudy (jdoe0001):

\(\large \sqrt{4-(-5)}\implies \sqrt{4-\times (-5)}\implies \sqrt{4-\times (-1\times 5)}\)

OpenStudy (isaiah.feynman):

The square root is defined, because of the following below \[\sqrt{a+b} \neq \sqrt{a}+\sqrt{b}\] So you MUST evaluate whatever is under the square root before taking the square root.

OpenStudy (dan815):

hahah

OpenStudy (dan815):

shes gonna get confused look at all this math and proofs for why negative * negative = positive

OpenStudy (anonymous):

im confused with this - Let h (x ) =sqrt 4 - x . Find h (0) +h (-5) because then it would be 9 and the sqrt of 9 is 3 and that is not one of my answers .........

OpenStudy (dan815):

is there an answer that says no

OpenStudy (anonymous):

but "undefined " is so I was pushing the -1

OpenStudy (anonymous):

h(0) = sqrt(4-0) = sqrt4 = 2 h(-5) = sqrt(4--5) = sqrt9 = 3 h(0) - h(-5) = 2 - 3 = -1

OpenStudy (anonymous):

if it's plus. h(0) + h(-5) = 5

OpenStudy (anonymous):

oh! got it i see now thanks!

OpenStudy (anonymous):

I can understand why you were getting confused. So was I. But Dan815 gave the best answer

OpenStudy (anonymous):

i totally didn't do the sqrt of 4

OpenStudy (anonymous):

haha